XioGonz:

Please help....with some trig.. Solve 2 sin(x)-2√3 cos(x)-√tan(x)+3=0 on[0,2π]

Ruthy:

Wich class is dis for

jhonyy9:

so like a first idea note sin(x) = a and cos(x) an b and write tan(x) = sin(x)/cos(x) so look what will get in this way

XioGonz:

I'm not sure I understand what you're trying to tell me.

jhonyy9:

use what i ve said above and get 2a - 2sqrt3(b) - sqrt (a/b) +3 = 0 so there inside radical is just tan (x) or tan(x)+3 ?

XioGonz:

tan(x)+3

jhonyy9:

ok than will be 2a +(2sqrt3)b - sqrt[(a/b) +3)] = 0 so but you need get an equation just with one variable - for this you can rewrite the cos(x) in function of sin(x) or viceversa the sin(x) in function of cos(x) using this equation sin^2 (x) +cos^2 (x) = 1

jhonyy9:

so do you understand the way ? you need get an equation with just one variable

XioGonz:

Ahhh I see.